OpenStudy (anonymous):

What is the solution of the following system? -3x-2y=-12 9x+6y=-9

OpenStudy (anonymous):

3(-3x-2y=-12) 9x+6y=-9 -9x-6y=-36 9x+6y=-9 if add down the x's, and y's cancel leaving you with 0=-45 which i can't recall but i think is indefinite solutions

OpenStudy (anonymous):

5sos

OpenStudy (anonymous):

It was actually no solutions but thank you for the help!

OpenStudy (anonymous):

i think calvin wood the asian cant sing...

OpenStudy (anonymous):

indefinite solution is consider no solutions.